OpenStudy (anonymous):

a rectangular field measures 50 yards by 100 yards. if it takes one quart of paint to paint a 15 foot by 30 foot section, how many gallons of paint will be needed to paint the entire field? (1 gallon=4 quarts)

OpenStudy (anonymous):

25 gallons

OpenStudy (anonymous):

work please

OpenStudy (anonymous):

50 * 100 = 5,000 sq. yards. 15 feet = 5 yards. 30 feet = 10 yards 5 * 10 = 50 sq yards. 5000/50 = 100 100/4 = 25